OpenStudy (anonymous):

If my overall grade is 77 currently and I have a quiz coming up that is worth 44% of the grade... and I get a 100 on it, what will my overall grade become? Or is that impossible to figure out.. lol

jimthompson5910 (jim_thompson5910):

it's hard to say because there are future assignments or tests that have yet to factor in but if we ignore that, then we can say that because this test is 44% of your grade, then everything else is 100-44 = 56% of your grade this allows us to say (56% of 77) + (44% of 100) = 0.56*77 + 0.44*100 = 87.12 so your grade would bump up to 87.12 ------------------------------------------------------- This would work if the test that's worth 44% was a final and no other work would need to be done. However, since other future work has yet to be accounted for, this means that the grade is probably going to be slightly less than 87.12

OpenStudy (anonymous):

now i def need to get a 100 on that quiz! lol thanks

jimthompson5910 (jim_thompson5910):

you're welcome I'm sure if you have a lot of assignments and tests coming up, then that provides more opportunity to bring your grade up but keep in mind that you have to do well on them of course